[ ESO ]
XSHOOTER Quality Control:
Known Problems

CAL | HC | refs | QC
QUALITY CONTROL
    HOME
XSHOOTER QC
Trending & QC1
known problems
Pipeline
Data Packages
Data Management
QC links:
Problems may also be related to the pipeline processing .

UVB  wrong acq. coord.  UVB ADC stuck  UVB pickup noise
VIS  VIS bias pattern  VIS remnance
NIR  NIR FF lamp  NIR remnance  K band vignetting




top Acquisition coordinates wrong for g_prime filter

From April 11 to May 14, 2011, the predicted coordinates for the slit center for the g_prime acquisition filter were wrong, which resulted in the placement of objects at the slit edge instead of the slit center and thus considerable flux loss.

top UVB Atmospheric Dispersion Compensator stuck

From July 18 to August 2, 2010, the Atmospheric Dispersion Compensator (ADC) of the UVB arm was stuck and did no longer adapt to changes in airmass or observing mode (SLIT vs. IFU). This resulted in strong vignetting of IFU data and in slit losses for SLIT data. Below you find a plot comparing observations of the same telluric standard stars with (red) and without (black) that problem (airmass noted besides date). The spectra were scaled to overlaps between 350nm and 400nm.

 

top Pickup noise in UVB data

The UVB CCD somtimes suffers from (horizontal) pickup noise, which is visible in low signal data and shows up in the rectified data as tilted lines. Below you see an example of raw data (rotated by 90 degrees) and merged rectified data. The example shown was observed in nodding mode, which creates the bright and dark region in the merged rectified data.

 

top Noise pattern and variable bias level in VIS data

Since early March 2010 the VIS CCD shows a time variable pattern and some pickup noise in the bias for the 400k readout mode. The amplitude of the pattern has been strongly reduced after March 20, but it is still present at a low level. Below you see an example of bias VIS raw frame (400k, low gain, 1x1, with cuts of 1025-1035 ADU and rotated clockwise by 90 degrees).

Additional work in October 2010 further reduced the pattern, albeit combined with a slight increase in read noise.

top Remnance in VIS data

Binned wavelength calibration data in the VIS arm tend to be highly saturated in some lines, which may leave remnance in data taken shortly after. The effects decay on the order of 20-60 minutes.

top Change in flux of NIR flat field lamp

Between November 1, 2009 and August 20, 2010 the NIR flat field lamp provided less flux in the K band than before and after. The effect of this on processed data can be clearly seen in the plot below, comparing a telluric standard star observed and processed before and after that change.

top Remnance in NIR data

NIR darks taken shortly after high flux observations (e.g. flat fields or flux standard stars in twilight) may show remnance in the top orders (raw data) or at right edge (processed data). Below you find an example of a 120s DARK, taken 30 minutes after a bright observations. The cuts are at 23 and 45 counts.

top Vignetting in K band

The upper part of the last orders in the K band shows 10% vignetting due to a small design error of the M8 baffle inside the NIR cryostat. This cancels out with flat fielding, but results in higher noise for this region.


 
[ESO][Index][Search][Help][News]